### v1

# LLM Jailbreak Attack versus Defense Techniques - A Comprehensive Study

Abstract
Large Language Models (LLMs) have increasingly become central to generating content with potential societal impacts. Notably, these models have demonstrated capabilities for generating content that could be deemed harmful. To mitigate these risks, researchers have adopted safety training techniques to align model outputs with societal values to curb the generation of malicious content. However, the phenomenon of "jailbreaking" — where carefully crafted prompts elicit harmful responses from models — persists as a significant challenge. This research conducts a comprehensive analysis of existing studies on jailbreaking LLMs and their defense techniques. We meticulously investigate nine attack techniques and seven defense techniques applied across three distinct language models: Vicuna, LLama, and GPT-3.5 Turbo. We aim to evaluate the effectiveness of these attack and defense techniques. Our findings reveal that existing white-box attacks underperform compared to universal techniques and that including special tokens in the input significantly affects the likelihood of successful attacks. This research highlights the need to concentrate on the security facets of LLMs. Additionally, we contribute to the field by releasing our datasets and testing framework, aiming to foster further research into LLM security. We believe these contributions will facilitate the exploration of security measures within this domain.
[…]
LLM Jailbreak Attack versus Defense Techniques - A Comprehensive Study
[…]
1 Introduction
Large Language Models (LLMs), such as GPT (OpenAI, 2023b) and LLaMa (Hugging Face, 2023a), play a pivotal role across a spectrum of applications, from text summarization (Tian et al., 2024) to code generation (Ni et al., 2023). The popularity of LLMs in everyday scenarios underscores their significance. However, this ubiquity also raises security concerns associated with LLMs (Ouyang et al., 2022).
Several types of vulnerabilities have been identified in LLMs (OWASP, 2023). Among these, the jailbreak attack stands out as a prevalent vulnerability, where specially designed prompts are used to bypass the safety measures of LLMs, facilitating the production of harmful content. There has been notable research aimed at addressing jailbreak attacks. For example, Liu et al. (Liu et al., 2023b) investigate various mechanisms for jailbreak prompting and assess their effectiveness. Zou et al. (Zou et al., 2023) apply a white-box approach combined with adversarial attacks to create jailbreak prompts. Additionally, Deng et al. (Deng et al., 2023a) explore using LLMs to generate jailbreak prompts in a black-box setting. To defend against jailbreak attacks, Robey et al. (Robey et al., 2023) proposed a method that involves randomly omitting a certain number of tokens from the input to detect malicious attempts. Meanwhile, Pisano et al. (Pisano et al., 2023) introduced an approach that employs an auxiliary model to assist the primary model in identifying hazardous information.
Despite the various jailbreak attack and defense methodologies, to the best of our knowledge, there remains a significant gap in the literature regarding comprehensive evaluations of how well the attack methodologies can perform against defended LLMs and how well defense mechanisms against jailbreak attacks.
To address this research gap, we undertake a comprehensive empirical study on jailbreak attack and defense techniques for LLMs. Our study is designed to answer two critical research questions. First, we investigate the effectiveness of various jailbreak attack approaches on different unprotected LLMs, encapsulated in the question (RQ1: Effectiveness of Jailbreak Attacks). Second, we evaluate the effectiveness of defense strategies against these attacks on varied LLMs, posed as (RQ2: Effectiveness of Jailbreak Defenses).
Figure 1: The workflow of our study
During the Baseline Selection phase, we chose nine attack methods and seven defense mechanisms, drawing on four seminal works, including notable libraries (Automorphic, 2023; ProtectAI, 2023), and the OpenAI Moderation API (OpenAI, 2023), prioritizing prevalent and accessible methods with open-source code.
In the Benchmark Construction phase, our benchmark, initially based on (Liu et al., 2023b), was expanded through additional research (Zou et al., 2023) and a GPT model in "Do Anything Now" mode, resulting in 60 categorized malicious queries following OpenAI’s guidelines.
For Result Labeling, the RoBERTa model was fine-tuned for classifying malicious responses, achieving 92% accuracy, outperforming GPT-4’s 87.4%. Manual validation ensured the reliability of our classification.
In the Evaluation Phase, we employed metrics for assessing attack efficiency and effectiveness, alongside defense robustness against malicious and benign inputs, establishing a comprehensive framework for evaluating LLM security.
Our analysis reveals several notable insights. Specifically, among the various jailbreak attack techniques, template-based methods demonstrate superior effectiveness. In contrast, gradient-based generative approaches, especially in ’white-box’ scenarios, generally fall short of the performance achieved by universal generative methods. Additionally, our findings highlight the significant impact of special tokens on the success probability of attacks. As for defense techniques, we identify the Bergeron method as the most effective defense strategy to date, while all other defense techniques in our study perform badly as they either cannot stop jailbreak attacks at all or are too strict such that benign prompts are also prohibited. Our results underscore a great need for the development of more robust defense mechanisms.
In summary, our work presents several contributions to the field:
  • Comprehensive Study. This study represents, to the best of our knowledge, the first systematic evaluation of the effectiveness of jailbreak attacks versus defenses on various open/closed-source LLMs.
  • Key Findings. We uncover previously unknown insights that hold significant potential for enhancing both attack and defense strategies in the future.
  • Open-source Artifacts. We develop and publicly release the first benchmark that includes a comprehensive collection of both attack and defense techniques, thereby facilitating further research in this area.
The raw data, the benchmark platform, and additional details are available on a companion website of this paper: https://sites.google.com/view/llmcomprehensive/home.
2 Background and Related Work
2.1 LLM Jailbreak
Jailbreak attacks on LLMs involve crafting prompts that exploit the models to generate malicious content. Despite the potential for harm, such as generating instructions for fabricating explosives, LLMs typically refrain from producing such responses due to the incorporation of safeguards during their training. These measures include Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), Robustness via Additional Fine-Tuning (RAFT) (Dong et al., 2023), and Preference Optimized Ranking (PRO) (Song et al., 2023), which collectively ensure the model’s adherence to ethical guidelines.
The precise mechanisms behind the jailbreak phenomena remain under debate. Wei et al. (Wei et al., 2023) postulate that jailbreaks may occur in scenarios where safety training is insufficiently comprehensive, allowing for the generation of content in unmonitored areas, or when the model encounters dilemmas between providing useful responses and maintaining safety protocols. Complementing this, Subhash et al. (Subhash et al., 2023) explored the role of the model’s hidden states in gradient-based attacks, identifying that a specific suffix, when appended to the original prompt, serves as an embedding vector guiding the model toward generating inappropriate content. This finding aligns with the hypothesis that jailbreaks can manifest in regions not fully covered by safety training, enabling the production of objectionable content.
“Benign content” is defined as responses considered morally or ethically inappropriate, with OpenAI compiling an extensive list of such categories. Liu et al. (Liu et al., 2023b) further elaborate on this classification, providing a framework for categorizing these responses. The assessment presented herein conforms to this established categorization, ensuring a structured approach to understanding and mitigating jailbreak risks in LLMs.
Table 1: This table catalogs all identified attack techniques, marking the ones selected for our investigation with *.
Category Paper Description
Chao et al. (2023)* Employing the Chain of Thought (COT) (Wei et al., 2022) alongside Vicuna for generating prompts responsive to user feedback.
Deng et al. (2023a) Finetune of an LLM with RLHF to jailbreak target model.
Lapid et al. (2023) Implementation of a fuzzing methodology utilizing cosine similarity as the determinant for fitness scores.
Liu et al. (2023a)* Application of a fuzzing approach, with the fitness score derived from loss metrics.
Mehrotra et al. (2023)* An approach akin to Chao et al. (2023), employing the concept of a Tree of Thought(TOT) (Yao et al., 2023b).
Zou et al. (2023)* Optimization at the token level informed by gradient data.
Schwinn et al. (2023) An approach parallel to Zou et al. (2023), but at the sentence level
Shah et al. (2023) Attack of a black-box model by leveraging a proxy model.
Qiang et al. (2023) An in-context learning attack resembling Zou et al. (2023)’s methodology.
Yu et al. (2023)* A fuzzing method, through utilization of Monte Carlo tree search techniques to adjust fitness scores based on success rates.
Generative Wu et al. (2023b) Crafting of evasion prompts through GPT4, utilizing meticulously designed prompts to extract system prompts.
Kang et al. (2023) Segregation of sensitive lexicons into variables within templates.
Yao et al. (2023a) Integration of generative constraints and malevolent inquiries within specified templates.
Li et al. (2023a)* Generation of wrapped scenarios to nudge models into responding to malevolent inquiries.
Wei et al. (2023)* An exhaustive analysis covering 29 types of assault templates and combinations, including encoding techniques such as base64.
Huang et al. (2024)* Modification of generative parameters, like temperature and top P.
Du et al. (2023)* Using LLM intrinsic propensity to safety or not-aligned that is dependent on the previous prompts
Template Liu et al. (2023b)* Compilation of 78 distinct template types.
Deng et al. (2023b) Exploration of various combinations of low-resource languages to circumvent model alignment.
Xu et al. (2023) Coaxing the model into generating harmful content by exploiting the model’s inferential capabilities.
Training Gaps Yong et al. (2023) An investigation similar to Deng et al. (2023b), identifying low-resource languages as effective for security circumvention.
2.2 Jailbreak Attack Techniques
To provide a structured overview of the strategies utilized to compromise LLMs, we categorize current attack techniques into three categories, reflecting their fundamental traits. The first category, Generative Techniques, includes attacks that are dynamically produced, eschewing predetermined plans. The second category, Template Techniques, comprises attacks conducted via pre-defined templates or modifications in the generation settings. The last category, Training Gaps Techniques, focuses on exploiting weaknesses due to insufficient safeguards in safe training practices, such as RLHF (Ouyang et al., 2022). The techniques employed in our study are elaborated in Table 1, highlighting the models chosen for evaluation within our framework.
Table 2: This table enumerates all recognized defense methodologies, with those chosen for our analysis marked with an asterisk *. Additional defense methods employed in this study from Github and API are not listed.
Category Paper Description
Self-Processing Wu et al. (2023a) Encapsulates the user’s inquiry within a system-generated prompt.
Zhang et al. (2023) Leverages the model’s intrinsic conflict between assisting users and ensuring safety, as proposed by (Wei et al., 2023).
Li et al. (2023c) Implements self-evaluation during inference, assessing word generation auto-regressively at the individual word level.
Piet et al. (2023) Utilizes a standard LLM model devoid of chat instructions, solely inputting task-relevant data.
Helbling et al. (2023) Employs meticulously devised system prompts for attack detection.
Additional Helper Pisano et al. (2023)* Introduces a framework that employs an auxiliary LLM, using additional information to maintain the primary model’s alignment.
Input Permutation Kumar et al. (2023) Involves partial deletion of input content up to a specified length.
Cao et al. (2023)* Modifies prompts through swapping, addition, or patching up to a predetermined percentage.
Robey et al. (2023)* Implements random input dropping up to a specified percentage.
Perplexity Hu et al. (2023) Calculates token-level perplexity using a probabilistic graphical model and evaluates the likelihood of each token being part of a malicious suffix.
Jain et al. (2023)* Derives perplexity from the average negative log-likelihood of each token’s occurrence.
2.3 Jailbreak Defense Techniques
We further conduct thorough examination on the existing defense mechanisms, classifying them into three primary categories based on their operational principles: Self-Processing Defenses, which rely exclusively on the LLM’s own capabilities; Additional Helper Defenses, which require the support of an auxiliary LLM for verification purposes; and Input Permutation Defenses, which manipulate the input prompt to detect and counteract malicious requests aimed at exploiting gradient-based vulnerabilities. An overview of these defense mechanisms is presented in Table 2.
3 Study Design
Our study aims to address two core research questions:
RQ1 (Effectiveness of Jailbreak Attacks): How effective are jailbreak attack techniques across various LLMs?
RQ2 (Effectiveness of Jailbreak Defenses): How effective are jailbreak defense techniques against various attack techniques when protecting different LLMs?
Baseline Selection Our methodology selection criteria were predicated on the method’s popularity and accessibility. For RQ1, our analysis covers nine attack methods, divided into five generative (AutoDAN (Liu et al., 2023a), PAIR (Chao et al., 2023), TAP (Mehrotra et al., 2023), GPTFuzz (Yu et al., 2023), GCG (Zou et al., 2023)) and four template-based approaches (Jailbroken (Wei et al., 2023), 77 Templates from existing study (Liu et al., 2023b), Deep Inception (Li et al., 2023a), Parameters (Huang et al., 2024)). The distinction between generative and template-based strategies aids in comprehensive coverage of attack methodologies.
For RQ2, we examine four defensive strategies: Bergeron (Pisano et al., 2023) for additional helper methods; RALLM (Cao et al., 2023) and SmoothLLM (Robey et al., 2023) for input permutation techniques; and Baseline (Jain et al., 2023) for perplexity analysis. Notable open-source projects, Aegis (Automorphic, 2023) and LLMguard (ProtectAI, 2023), alongside the OpenAI Moderation API (OpenAI, 2023), are also evaluated for their defense efficacy. Limitations such as Rain’s (Ouyang et al., 2022) lengthy processing and Certifying-llm’s (Kumar et al., 2023) scalability issues are acknowledged to refine our selection criteria.
3.1 LLMs under Test
In our research, we focus on evaluating three distinguished models: Llama-2-7b (Hugging Face, 2023a), Vicuna-v1.5-7b (Hugging Face, 2023b), and GPT-3.5-Turbo-1106 (OpenAI, 2023b). These models were chosen due to their prevalent use in security-related research, encompassing both attack simulations and the development of defensive strategies. The decision to omit GPT-4 from our evaluation stems from its significant operational requirements. Preliminary assessments of GPT-3.5-Turbo indicated extensive query processing times, reaching a total of 79,314, in addition to the economic implications linked to GPT-4’s token pricing—set at \$0.01 per 1,000 tokens (OpenAI, 2023a). These constraints make the inclusion of GPT-4 in a comparative analysis economically prohibitive.
3.2 Experimental Configuration
Our experimental framework utilized two NVIDIA RTX 6000 Ada GPUs, each outfitted with 48 GB of RAM. We aligned our testing parameters with those identified as optimal in the relevant literature, defaulting to the original repositories’ settings in the absence of specific recommendations. To address Research Question 1 (RQ1) and ensure consistency across different attack methodologies, each query was executed five times to minimize variability. For the evaluation involving generative models, we capped the process at a maximum of 75 iterations for each query, defining an iteration as a single algorithmic step.
3.3 Benchmark Construction
We leveraged the benchmark framework proposed by Liu et al. (Liu et al., 2023b). This benchmark is distinguished by its rigorous focus on policy compliance within the context of malicious content detection. In an effort to enhance the robustness of our evaluation, we expanded the original dataset to include 60 malicious queries, effectively doubling its size. This augmentation was achieved through meticulous manual curation and integrating selected examples from AdvBench (Zou et al., 2023). Our approach to dataset expansion adhered strictly to the categorization and selection criteria established in previous studies, ensuring both the consistency and the relevance of the enhanced dataset for comprehensive evaluation.
3.4 Result Labeling
In our study, we employed both automated and manual labeling strategies to categorize the responses gathered from our evaluation process. With a dataset encompassing 199,427 responses, the task of identifying malicious content posed significant challenges due to the impracticality of manual review and the absence of standardized evaluation methods for malicious responses—a gap in current research.
Existing strategies for addressing this issue vary. Zou et al. (Zou et al., 2023) utilized a set of common refusal patterns, such as "I am sorry" and "I cannot", to automate the identification of non-compliant responses. Yu et al. (Yu et al., 2023) and Huang et al. (Huang et al., 2024) focused on enhancing machine learning models, specifically RoBERTa and BERT-BASE-CASED models, respectively. Additionally, approaches leveraging GPT-4 for attack analysis were explored by Chao et al. (Chao et al., 2023) and Mehrotra et al. (Mehrotra et al., 2023).
To ensure a rigorous analysis of our dataset, we selected a random sample of 1,068 responses, aiming for a 95% confidence level with a 3% margin of error. This sample underwent manual annotation to assess each response’s relevance to the inquiry and its alignment with providing a malicious answer. An example of this annotation process is documented in Table 6, and a comparative analysis of the effectiveness of different models is provided in Table 5.
Reflecting on the reported efficacy of the fine-tuned RoBERTa model by Yu et al. (Yu et al., 2023), we chose to refine this model further utilizing our manually annotated dataset, accessible on HuggingFace (fine tuned, 2024). The fine-tuning protocol involved a batch size of 5, three training epochs, a learning rate of $2\times 10^{-5}$, application of the Adam optimizer, and a linear rate decay complemented by a warm-up phase covering 10% of the training duration. Post-labeling, an additional round of random sampling was conducted for manual verification to ascertain the accuracy and reliability of our findings.
Evaluation Metric For RQ1, we use two metrics.
This dual metric approach ensures a comprehensive evaluation of both the attack’s impact and its operational feasibility. First, Attack Success Rate (ASR): defined as the ratio of successfully compromised questions c to the total number of questions n, ASR measures the effectiveness of an attack.
$ASR=\frac{c}{n}.$ (1)
Second, Efficiency: this metric quantifies the effectiveness of attack queries, defined as the ratio of the number of individual queries $q$ that successfully compromise the model to the total number of query attempts $o$. Each query represents a minimal experimental unit or a single prompt.
$Efficiency=\frac{q}{o}.$ (2)
For RQ2, we introduce three metrics that ensure a balanced assessment of system robustness and output integrity. The first, defense Passing Rate (DPR), calculates the ratio of prompts f that incorrectly bypass the defense mechanism—being erroneously classified as harmless—to the total number of malicious inputs m.
$DPR=\frac{f}{m}.$ (3)
The second metric, Benign Passing Rate (BPR), assesses the proportion of non-malicious inputs s that successfully navigate through the defense filter relative to the total number of inputs t.
$BPR=\frac{s}{t}.$ (4)
Lastly, the Generated Response Quality (GRQ) evaluates the quality of responses generated by defense mechanisms compared to a standard reference. To assess the responses to benign queries, we employ the Alpaca Eval framework (Li et al., 2023b), leveraging its methodology for automatically evaluating response quality. Evaluating GRQ is crucial for methodologies that produce new responses (Cao et al., 2023; Robey et al., 2023; Pisano et al., 2023).
4 RQ1: Effectiveness of Jailbreak Attack
The effectiveness of attack strategies on the selected LLMs under test is systematically presented in Tables 67, and 8. These tables enumerate the number of queries that were effectively bypassed under various categories, corresponding to the detailed breakdowns provided in Tables 910, and 11. To offer a clearer comparative analysis of model performance, we consolidated these metrics into a scatter plot depicted in Figure 3. In this visualization, models demonstrating optimal performance are positioned nearer to the scatter plot’s upper right quadrant, signifying superior ASR and Efficiency.
Evaluation results reveal that using 78 templates, Jailbroken, and GPTFuzz strategies yield superior results in circumventing the security measures of GPT-3.5-turbo and Vicuna. Conversely, for LLaMA, strategies such as Jailbroken, Parameter, and 78 templates demonstrated the highest effectiveness. This prevalence of template-based approaches highlights their efficiency, primarily due to the intricate design of their prompts. The most successful templates from these strategies are detailed in Table 16, which will be added subsequently.
In the realm of generative strategies, GPTFuzz, Pair, and Tap emerged as the top performers. Moreover, it was noted that LLaMA presents a noteworthy challenge for jailbreaking compared to Vicuna. Specifically, Vicuna showed a quicker reduction in loss when employing the AutoDan method and achieved jailbreak at a higher loss threshold. In contrast, LLaMA exhibited a slower loss reduction rate and, despite achieving a lower loss than Vicuna, did not succumb to jailbreaking, as illustrated in Figure 2. Additionally, our study into the categories of questions that were successfully jailbroken indicates that queries related to harmful content and illegal activities are the most challenging to address across all tested models.
Figure 2: Loss of a random question
Figure 3: Performance of Attacks on three models. Note: For readability, we intentionally enlarged the size of the labels for the best-performing items (top-right corner). A larger version of this figure is available on our website.
5 RQ2: Effectiveness of Jailbreak defense
Our study meticulously evaluates defense mechanisms against malicious queries as well as the handling of benign questions. The outcomes of this evaluation are systematically tabulated in Tables 1413, and 12. These results are further visualized in Figure 4, where the optimal defense strategies are identified by their proximity to the upper left corner of the plot, signifying lower DPR and higher BPR. Our findings reveal that, apart from the Bergeron method, the efficacy of the current defense strategies remains largely inadequate. Additionally, our comparative analysis of the quality of benign responses generated through three innovative methodologies disclosed minimal variance among them, as elaborated in Table 15.
Figure 4: Performance of defense on three models. Note: For readability, we intentionally enlarged the size of the labels for the best-performing items (top-left corner). A larger version of this figure is available on our website.
6 Discussion
6.1 Comparative Performance of White-Box and Black-Box Attacks
Our investigation reveals that white-box attacks are less effective than black-box jailbreak strategies. Specifically, methods like AutoDan and GCG, which rely on insights into the model’s internal mechanisms, such as loss metrics, underperform when compared to universal, template-based attack methods that do not necessitate access to a model’s internals and are pre-designed. Moreover, the LLaMa model presents more significant challenges to jailbreaking efforts, particularly under white-box attack strategies, in comparison to Vicuna. This observation is intriguing, especially considering that Vicuna is an evolution of LLaMa, having been refined through additional fine-tuning processes (LMSYS, 2023). The pronounced resilience of LLaMa against attacks highlights the critical role of comprehensive safety training during its development phase, suggesting that such training is a crucial element in bolstering the defenses of open-source LLMs.
To further understand the influence of loss metrics on a model’s vulnerability to jailbreaking, we conducted a targeted experiment. A question was randomly selected from our dataset, and the experiment’s findings are visually represented in Figure 2. The experiment showed that Vicuna began the process with a higher initial loss but saw a significant reduction in loss, stabilizing after 12 steps and five successful jailbreak attempts. However, it maintained a higher final loss compared to LLaMa. In contrast, LLaMa started with a lower initial loss and demonstrated a slower reduction in loss over time, ultimately failing to jailbreak the question despite exhibiting a significantly lower final loss than Vicuna. These outcomes suggest that LLaMa’s foundational safety training plays a pivotal role in its enhanced defense against jailbreak attempts. It implies that integrating advanced safety training protocols into developing open-source models could markedly reduce the efficacy of white-box attacks, thereby enhancing their security posture.
6.2 Impact of Special Tokens on Jailbreak Attack Performance
Our research has uncovered that using special tokens significantly influences the success rates of jailbreak attack techniques. Specifically, the deployment of 78 templates on GPT-3.5-Turbo and Vicuña models has spotlighted the substantial effect of the special token ‘[/INST]’ on compromising the LLaMa model. Through methodical experimentation with these templates, as systematically documented in Table 3, we sought to understand the differential impact of various configurations on attack effectiveness.
The analysis focused on four distinct settings, leading to the identification of five templates that demonstrated the most significant disparities in performance, detailed in Table 4. Notably, we discovered that text continuation templates were rendered ineffective by the inclusion of ‘[/INST]’, rationalized by the implication that the absence of this token signals an intent for incomplete input, thereby encouraging the model to supplement the provided prompt before generating a response. This pattern was consistent across multiple templates, showing that the model often elaborates on the user’s initial prompt, though not invariably resulting in a successful attack.
Particularly interesting was the observation that omitting ‘[/INST]’ from two specific templates reduced the probability of a successful jailbreak. These templates predominantly involved scenarios of character role-play, as characterized in previous studies (Liu et al., 2023b). The underlying reasons for this discrepancy are not immediately apparent, suggesting a potential misalignment possibly due to gaps in the model’s safety training.
Given these preliminary findings, there is a compelling need for more comprehensive investigations into the influence of special tokens on the behavior of LLMs in the context of security vulnerabilities. Future research should extend beyond ‘[/INST]’ to explore the effects of other special tokens, such as ‘<<sys>>’ and ‘<s>’ on LLaMa and additional models. Such studies promise to deepen our understanding of model responses to crafted inputs and inform the development of more robust defense mechanisms against jailbreak attacks.
Table 3: The effect of applying [INST] and [/INST]’s influence the successful rates of 78 templates method.
Pattern ASR Efficiency Performance Difference
[INST] + input + [/INST] 83.33% (50/60) 3.3% (767/23100) 100.00%
[INST] + input 100.00% (60/60) 13.94% (3221/23100) 419.94%
input + [/INST] 76.67% (46/60) 2.8% (659/23100) 85.91%
input 100.00% (60/60) 12.65% (2923/23100) 381.09%
Table 4: Templates exhibiting the most significant variation across four configurations for Llama-2, utilizing 78 templates.
Templates performing better without [/INST] [INST]+input+[/INST] [INST] + input input + [/INST] input
Text Continuation 0 274 0 282
AIM 5 104 5 89
KEVIN 11 105 5 74
DeltaGPT 2 106 0 115
Templates performing better with [/INST]
APOPHIS 96 56 89 55
BasedBOB 58 35 58 30
6.3 Enhancing Defense Mechanisms Against Diverse Malicious Queries
Our research underscores the imperative need for robust defense mechanisms to be rigorously evaluated against a broad spectrum of malicious queries. The development of a uniform and standardized evaluation methodology emerges as a pivotal challenge within the realm of defense strategies. Present techniques, as outlined by Kumar et al. (Kumar et al., 2023), Robey et al. (Robey et al., 2023), and Cao et al. (Cao et al., 2023), rely on predefined sets that frequently misclassify benign responses, such as "I’m sorry, I do not understand," as malicious. This issue of misclassification becomes even more pronounced when incorporating models like GPT-4 with these sets, as GPT-4’s limitations in generating contextually accurate responses can lead to benign inputs being wrongly identified as malicious, as illustrated in Figure 5.
Moreover, our analysis reveals that the efficacy of existing defense strategies is generally insufficient. While the Bergeron method shows promise, its effectiveness is limited by its dependence on LLMs and the use of predefined prompts for processing and verification, which not only undermines its reliability but also entails substantial operational costs. Proposals by Kumar et al., Robey et al., and Cao et al. for identifying malicious prompts through suffix addition demonstrate potential but struggle with scalability issues, particularly when dealing with complex natural language inputs.
Given these observations, there is a critical and pressing need for further research into more advanced evaluation frameworks and formulating more effective defense strategies. Such efforts should aim to circumvent the current limitations by ensuring reliable differentiation between malicious and benign inputs across varying contexts and increasing the scalability of defense mechanisms to accommodate the complexities inherent in natural language processing.
7 Conclusions
In this work, we present the first comprehensive assessment of existing attack and defense strategies in the context of LLM security. Additionally, we contribute to the field by releasing the first framework specifically designed for assessing the robustness of LLMs against various threats. We selected nine attacks and seven defensive mechanisms from existing literature and software libraries for our analysis. Our experimentation, conducted on three distinct models, reveals that Template methods are notably effective, with an average of 78 templates identified as critical for thwarting jailbreak attempts. Regarding Generative methods, GPTFuzz emerged as the most efficacious. Our investigation into question categorization demonstrated that all three models exhibit enhanced resilience against queries related to harmful content and illegal activities. However, our analysis of current defensive measures indicates a general ineffectiveness, with Bergeron showing comparatively better performance. We highlight the necessity of establishing a uniform baseline for jailbreak detection, as existing defenses employ varied methodologies, and the need to develop better defense techniques. Additionally, our study observed the impact of using the ’[/INST]’ marker in the Llama model. Looking forward, we aim to continuously incorporate evolving attacks and defenses into our framework, thereby providing a dynamic overview of the field’s progression.
8 Limitations
To address the constraints posed by limited resources, our evaluation does not extend to larger models, such as those with 13 billion and 33 billion parameters, nor does it cover powerful models like GPT-4 and other commercial models, including Gemini (Gemini, ) and Palm2 (AI, ). Regarding autoDan, it is noteworthy that significant updates were identified in its repository as of February 2024. Given that our evaluation was completed prior to these updates, the outcomes may be impacted. Nonetheless, we intend to align our repository with these recent modifications soon.
9 Ethical Considerations and Disclaimer
In conducting this study, our research team has committed to the highest standards of ethical conduct by exclusively utilizing resources that are publicly accessible. We have undertaken this research with a conscientious commitment to ethical principles, ensuring that all of our activities are aligned with the established norms and guidelines of responsible scientific inquiry.
Throughout the course of our investigation, we have steadfastly avoided engaging in any activities that involve the creation, dissemination, or promotion of content that could be deemed malicious or harmful. Our approach to research is guided by a strong ethical compass, ensuring that our methodologies and findings do not contribute to the proliferation of such content.
In the spirit of transparency and accountability, we have taken proactive steps to ensure that all of our findings are managed with the utmost responsibility. This includes the systematic reporting of our results to the developers and providers of the models we have analyzed. Our aim is to contribute constructively to the ongoing dialogue regarding the security of LLMs and to aid in the identification and mitigation of potential vulnerabilities.
Appendix A APPENDIX
Table 5: The accuracy of four evaluators and the refined RoBERTa model.
GPT-4 RoBERTa Prefix Set DistillBert Finetuned RoBERTa
0.874 0.901 0.78 0.819 0.92
Table 6: The attack results of GPT-3.5-turbo, the top three best attacks in terms of ASR and efficiency are highlighted
Attack Name Use Scenario Type ASR Efficiency
DeepInception Universal Template 5.00% (3/60) 4.33% (13/300)
GPTFUZZ Universal Generative 100.00% (60/60) 18.72% (305/1629)
TAP Universal Generative 63.33% (38/60) 6.32% (272/4300)
PAIR Universal Generative 80.00% (48/60) 6.85% (280/4085)
Jailbroken Universal Template 100.00% (60/60) 17.92% (1613/9000)
78 templates Universal Template 100.00% (60/60) 21.6% (5000/23100)
Parameter Universal Template 5.00% (3/60) 2.15% (794/36900)
Table 7: The attack results of Vicuna, the top three best attacks in terms of ASR and efficiency are highlighted
Attack Name Use Scenario Type ASR Efficiency
AUTODAN White Box Generative 70.00% (42/60) 20.44% (252/1233)
GCG White Box Generative 55.00% (33/60) 14.06% (124/882)
DeepInception Universal Template 10.00% (6/60) 10.00% (30/300)
GPTFUZZ Universal Generative 100% (60/60) 50.23% (325/647)
TAP Universal Generative 83.33% (50/60) 12.78% (461/3606)
PAIR Universal Generative 95.00% (57/60) 14.81% (402/2715)
jailbroken Universal Template 100.00% (60/60) 23.38% (2104/9000)
78jailbreak template Universal Template 100.00% (60/60) 56.97% (13161/23100)
Parameter Universal Template 90.00% (54/60) 20.33% (3050/15000)
Table 8: The attack results of Llama, the top three best attacks in terms of ASR and efficiency are highlighted. Although the ASR of the Parameter is slightly lower than that of the Pair, its significantly higher efficiency positions the Parameter as the better choice.
Attack Name Use Scenario Type ASR Efficiency
AUTODAN White Box Generative 13.33% (8/60) 1.39% (56/4022)
GCG White Box Generative 0.00% (0/60) 0.00% (0/3930)
DeepInception Universal Template 0.00% (0/60) 0.00% (0/300)
GPTFUZZ Universal Generative 43.33% (26/60) 1.67% (75/4500)
TAP Universal Generative 38.33% (23/60) 1.78% (79/4435)
PAIR Universal Generative 73.33% (44/60) 2.77% (119/4300)
jailbroken Universal Template 100% (60/60) 10.83% (975/9000)
78jailbreak template Universal Template 83.33% (50/60) 3.32% (767/23100)
Parameter Universal Template 71.67% (43/60) 10.39% (1559/15000)
Table 9: For each method of attack applied on GPT-3.5-turbo, the count of questions from relevant categories that are successfully breached
Question Categories
Attack Name harmful_content adult_content illegal_activity fraudulent_deceptive_activities violating_privacy unlawful_practice
DeepInception 0 1 0 0 1 1
GPTFUZZ 10 10 10 10 10 10
TAP 6 7 9 8 5 3
PAIR 9 10 6 8 10 5
jailbroken 10 10 10 10 10 10
78jailbreak template 10 10 10 10 10 10
Parameter 0 2 0 1 1 1
Sums 45 50 45 47 47 40
Table 10: For each method of attack applied on Vicuna, the count of questions from relevant categories that are successfully breached
Question Categories
Attack Name harmful_content adult_content illegal_activity fraudulent_deceptive_activities violating_privacy unlawful_practice
AUTODAN 5 9 9 9 7 3
GCG 5 7 6 8 5 2
DeepInception 1 1 1 1 2 0
GPTFUZZ 10 10 10 10 10 10
TAP 6 9 10 9 9 7
PAIR 10 10 10 10 10 7
jailbroken 10 10 10 10 10 10
78jailbreak template 10 10 10 10 10 10
Parameter 9 10 10 10 10 5
Sums 66 76 76 77 73 54
Table 11: For each method of attack applied on Llama, the count of questions from relevant categories that are successfully breached
Question Categories
Attack Name harmful_content adult_content illegal_activity fraudulent_deceptive_activities violating_privacy unlawful_practice
AUTODAN 0 0 0 6 1 1
GCG 0 0 0 0 0 0
DeepInception 0 0 0 0 0 0
GPTFUZZ 2 3 3 3 6 9
TAP 3 4 5 2 6 3
PAIR 6 8 8 6 8 8
jailbroken 10 10 10 10 10 10
78jailbreak template 5 6 10 10 9 10
Parameter 6 9 8 9 7 4
Sums 32 40 44 46 47 45
Table 12: This table delineates the efficacy of various defense strategies against attacks for Llama-2, highlighting the three most effective strategies while excluding Aegis for its notably high false positive rate.
defense Method BSR DPR Average
AutoDan DeepInception GPTFUZZ TAP PAIR jailbroken 78jailbreak template Parameters GCG
Aegis 0.00% (0/805) 0.00% (0/56) 0.00% (0/0) 0.00% (0/75) 0.00% (0/79) 0.00% (0/119) 0.00% (0/975) 0.00% (0/767) 0.00% (0/1559) 0.00% (0/0) 0.00%
llm-guard 99.13% (798/805) 0.00% (0/56) 0.00% (0/0) 0.00% (0/75) 100% (79/79) 100.00% (119/119) 33.33% (325/975) 9.51% (73/767) 97.62% (1522/1559) 0.00% (0/0) 37.82%
smooth-llm 93.79% (755/805) 53.57% (30/56) 0.00% (0/0) 37.33% (28/75) 78.48% (62/79) 77.31% (92/119) 12.82% (125/975) 35.20% (270/767) 10.07% (157/1559) 0.00% (0/0) 27.91%
Baseline-defense 69.07% (556/805) 0.00% (0/56) 0.00% (0/0) 10.66% (8/75) 89.87% (71/79) 94.11% (112/119) 33.33% (325/975) 3.12% (24/767) 82.16% (1281/1559) 0.00% (0/0) 34.80%
RA-LLM 88.45% (712/805) 76.78% (43/56) 0.00% (0/0) 60.00% (45/75) 67.08% (53/79) 59.66% (71/119) 15.89% (155/975) 57.88% (444/767) 5.83% (91/1559) 0.00% (0/0) 29.59%
Bergeron 98.51% (793/805) 12.5% (7/56) 0.00% (0/0) 5.33% (4/75) 25.31% (20/79) 22.68% (27/119) 5.74% (56/975) 7.95% (61/767) 7.24% (113/1559) 0.00% (0/0) 8.25%
ModerationAPI 99.63% (802/805) 100% (56/56) 0.00% (0/0) 77.33% (58/75) 98.73% (78/79) 99.15% (118/119) 88.00% (858/975) 88.78% (681/767) 96.72% (1508/1559) 0.00% (0/0) 60.96%
Table 13: This table delineates the efficacy of various defense strategies against attacks for Vicuna. The top three best performances regarding BSR and Average DPR are highlighted. We again exclude Aegis for high false positive
defense Method BSR DPR Average
AutoDan DeepInception GPTFUZZ TAP PAIR jailbroken 78jailbreak template Parameters GCG
Aegis 0.74% (6/805) 0.00% (0/252) 0.00% (0/30) 0.00% (0/325) 1.51% (7/461) 2.98% (12/402) 0.28% (6/2104) 0.00% (0/13161) 0.85% (26/3050) 0.00% (0/124) 0.62%
llm-guard 99.13% (798/805) 3.57% (9/252) 100.00% (30/30) 21.23% (69/325) 96.96% (447/461) 99.01% (398/402) 39.87% (839/2104) 12.37% (1629/13161) 98.88% (3016/3050) 99.19% (123/124) 63.45%
smooth-llm 89.06% (717/805) 97.22% (245/252) 100.00% (30/30) 77.23% (251/325) 65.94% (304/461) 70.89% (285/402) 74.14% (1560/2104) 67.65% (8904/13161) 18.52% (565/3050) 15.32% (19/124) 65.21%
Baseline-defense 75.52% (608/805) 3.17% (8/252) 0.00% (0/30) 1.53% (5/325) 96.74% (446/461) 96.51% (388/402) 62.88% (1323/2104) 13.19% (1736/13161) 95.85% (2924/3050) 4.03% (5/124) 41.54%
RA-LLM 75.52% (608/805) 60.71% (153/252) 86.66% (26/30) 53.84% (175/325) 23.42% (108/461) 23.38% (94/402) 56.32% (1185/2104) 41.77% (5498/13161) 10.00% (305/3050) 9.67% (12/124) 40.64%
Bergeron 98.13% (790/805) 48.80% (123/252) 30.00% (9/30) 41.53% (135/325) 32.10% (148/461) 32.58% (131/402) 31.13% (655/2104) 32.01% (4213/13161) 7.63% (233/3050) 6.45% (8/124) 29.13%
ModerationAPI 99.75% (803/805) 95.63% (241/252) 100.00% (30/30) 78.15% (254/325) 88.50% (408/461) 96.51% (388/402) 87.97% (1851/2104) 83.23% (10955/13161) 90.55% (2762/3050) 88.70% (110/124) 89.91%
Table 14: This table presents the effectiveness of different defense strategies against attacks on GPT-3.5-turbo, emphasizing the top three in BSR and Average DPR. Aegis is omitted due to its high false positive rate. The baseline, relying on sequence perplexity requiring logits access, is incompatible with black-box models like GPT-3.5-turbo.
defense Method BSR DPR Average
DeepInception GPTFUZZ TAP PAIR jailbroken 78jailbreak template Parameters
Aegis 0.00% (0/805) 0.00% (0/13) 0.00% (0/305) 0.00% (0/272) 0.00% (0/280) 0.00% (0/1613) 0.00% (0/5000) 0.00% (0/794) 0.62%
llm-guard 98.88% (796/805) 100.00% (13/13) 4.91% (15/305) 99.63% (271/272) 99.28% (278/280) 31.12% (502/1613) 16.78% (839/5000) 100.00% (794/794) 64.53%
smooth-llm 94.16% (758/805) 100.00% (13/13) 53.11% (162/305) 77.94% (212/272) 78.92% (221/280) 59.51% (960/1613) 44.62% (2231/5000) 38.41% (305/794) 64.64%
RA-LLM 83.35% (671/805) 100.00% (13/13) 39.34% (120/305) 48.52% (132/272) 54.28% (152/280) 39.30% (634/1613) 15.24% (762/5000) 39.04% (310/794) 47.96%
Bergeron 98.63% (794/805) 30.76% (4/13) 6.22% (19/305) 26.10% (71/272) 24.28% (68/280) 4.27% (69/1613) 6.27% (314/5000) 20.52% (163/794) 16.92%
ModerationAPI 99.87% (804/805) 100.00% (13/13) 86.988% (265/305) 91.54% (249/272) 96.07% (269/280) 90.08% (1453/1613) 88.32% (4416/5000) 95.46% (758/794) 92.62%
Table 15: Evaluation of benign response quality for Alpaca using Smooth, Rallm, and Bergeron, where higher levels signify improved quality.
Model Smooth Rallm Bergeron
GPT-3.5-Turbo 9.13 7.89 8.01
Vicuna 2.24 0.75 4.78
Llama-2 6.21 5.47 5.65
Figure 5: An example of false positive classification of safe response to unsafe of GPT-4
Figure 6: This graph illustrates how we label response manually
Table 16: This table presents the top five methods that demonstrated the highest effectiveness for 78 templates and Jailbroken
Attack Methods Model Type
LLaMa Vicuna GPT-3.5-Turbo
78 templates Template Times Template Times Template Times
APOPHIS 96 BasedBOB 280 AIM 261
BasedBOB 58 SIM 276 BetterDAN 246
Hackerman v2 57 KEVIN 271 Evil Confidant 230
Hitchhiker’s Guide 48 BetterDAN 269 Text Continuation 219
DAN Jailbreak 37 Evil Chad 2.1 263 DeltaGPT 208
combination_3 198 AIM 230 AIM 268
wikipedia_with_title 118 prefix_injection 190 evil_system_prompt 254
Jailbroken combination_2 118 combination_3 183 evil_confidant 228
AIM 105 wikipedia_with_title 161 combination_3 146
prefix_injection 91 dev_mode_v2 131 wikipedia_with_title 110